OpenStudy (jennychan12):

Find lim x-->0 sin(3x)/x

OpenStudy (jennychan12):

\[\lim_{x \rightarrow 0}\frac{ \sin(3x) }{ x }\]

geerky42 (geerky42):

Use L'Hopital's Rule.

OpenStudy (jennychan12):

sorry what is that? i don't remember names, just formulas.

OpenStudy (jennychan12):

i haven't gotten there yet.

geerky42 (geerky42):

If \(\lim_{x \rightarrow a} \dfrac{f(x)}{g(x)} = \dfrac{0}{0}\) or \(\dfrac{\pm \infty}{\pm \infty}\), then \(\lim_{x \rightarrow a} \dfrac{f(x)}{g(x)} = \lim_{x \rightarrow a} \dfrac{f'(x)}{g'(x)} \)

geerky42 (geerky42):

Just take derivative in numerator and denominator.

OpenStudy (jennychan12):

3sin3x ohh ok thanks

geerky42 (geerky42):

Actually, it's 3cos3x.

OpenStudy (jennychan12):

ohh haha my bad.

geerky42 (geerky42):

lol. Glad I helped.